doctor order 2ml ampule of 50 mg.Available is 60mg/3ml.How much ml nurse will give?
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

      Formula = Desired dose/Dose in hand *Quantity.

       Here

           Desired dose(D) = 50 mg.

           Dose in hand (H) = 60 mg.

                Quantity(Q) = 3 ml.

   As per above formula

          Volume(ml) of drug required = D/H * Q

                                                        = 50/60 * 3

                                                        = 0.83 * 3

                                                        = 2.5 ml.